Suppose the company will not pay any dividends in Years 1 and 2. Suppose that the company pays dividend of $1 in Year 3 and after that the dividends will grow at 20% for the next two years. After that the dividends will grow at a constant rate of 5% forever. If the required rate of return is 10%, compute today's price of the stock.
Answer #1

Cash Flow in year 3, CF3= $1

PV of CF3= 1/(1+0.1)^3=$0.75

CF4= $1*(1+20%)= $1.2

PV of CF3= 1.2/(1+0.1)^4=$0.82

CF5= $1.2*(1+20%)= $1.44

PV of CF3= 1.44/(1+0.1)^5=$0.89

After that CFs will grow at 5%. So, PV of all future CFs starting at year 6= $1.44*(1.05)/[(0.1-0.05)*(1+0.1)^6]=$17.06

Sum of all PVs= $19.53= Price of stock today